arooj is 20 years than her mother. after 6 years the sum of their ages will be 50 years. find their present ages?
Answer by rm29924(97)

Let Arooj's mother's current age = X
Arooj's current age is X-20
Arooj's mother's age in 6 years= X+6
Arooj's age in 6 years= X-20+6
after 6 years the sum of their ages will be 50 years so we add there ages in 6 years and set it equal to 50
X+6+X-20+6=50
combine like terms
2X-8=50
add 8 to both sides
2X=58
divide both sides by 2 to get the variable X by itself
X=29
Arooj's mother is 29 years old
Arooj is X-20=29-20=9 Arooj is 9 years old